Sorry, this was a section that I was not able to cover due to schedule changes. Thank You!

The problem: Rhonda mixes some 35% acid solution with some 70% acid solution to make 100mL of a 42% acid solution. How much of the 35% acid solution did she use?
Again, Thank You for your time,
A student

Answer by oberobic(2304)

With mixture problems, you need to determine how much 'pure' stuff you have or need.
.
She needs to make 100 mL @ 42% = 42 mL of pure acid.
.
The two amounts will total 100 mL, so she adds
x mL of 35% acid with
100-x mL of 70% acid
.
.35x + .7(100-x) = .42(100)
.
multiply both sides by 100 to eliminate decimals
.
35x + 70(100-x) = 42(100)
.
35x + 7000 -70x = 4200
.
-35x = -2800
.
x = -2800/-35 = 80 mL of 35% acid
.
100-x = 20 mL of 70% acid
.
check how much pure acid you have to be sure
.
.35*80 = 28 mL
.70*20 = 14 mL
28+14 = 42 mL
correct
.
Answer: Mix 80 mL of 35% acid + 20 mL of 70% acid to produce 100 mL of 42% acid.
.
Done.